OpenStudy (anonymous):

The are of the circle is 60piin^2 . find the circumference

OpenStudy (anonymous):

pi

OpenStudy (akashdeepdeb):

pi*r*r = 60*pi r = sqrt{60} So put this value in Circumference formula. :)

OpenStudy (anonymous):

square root of 60 does not simplify to a whole number so it would be 2sqrt15?